OpenStudy (anonymous):

Find the exact value of the volume of a can of oil with a diameter 10 centimeters and height of 12 centimeters. Include correct units with your solution. Use "pi" in place of the pi symbol.

OpenStudy (amistre64):

the volume of a cylinder = base area * height

OpenStudy (amistre64):

the base area is a circle; so use pi r^2 to determines it value

OpenStudy (amistre64):

diam = 10; 2 = d/2 so... 5 in this case 5^2 pi * 12 = 25*12 pi = 300 pi? 12 quarters = 3 dollars right?

OpenStudy (anonymous):

ok. thank you.